Find the area, take the elements of the area perpendicular to the x-axis. x²-y+1=0; x-y+1=0.

1
Expert's answer
2022-05-19T18:48:24-0400
"x^2-y+1=0=>y=x^2+1"

"x-y+1=0=>y=x+1"

"x^2+1=x+1"

"x_1=0, x_2=1"


"y(0)=0+1=1"

"y(1)=1+1=2"


"x^2-y+1=0, x\ge0"

"x=\sqrt{y-1}"

"x-y+1=0=>x=y-1"

"A=\displaystyle\int_{0}^{1}(x+1(-x^2+1))dx""=\bigg[\dfrac{x^2}{2}-\dfrac{x^3}{3}\bigg]\begin{matrix}
 1 \\
 0
\end{matrix}""=\dfrac{1^2}{2}-\dfrac{1^3}{3}-0=\dfrac{1}{6}({units}^2)"

"-(\dfrac{2(1-1)^{3/2}}{3}-\dfrac{(1)^2}{2}+1)=\dfrac{1}{6}({units}^2)"

"A=\displaystyle\int_{1}^{2}(\sqrt{y-1}-(y-1))dy""=\bigg[\dfrac{2(y-1)^{3/2}}{3}-\dfrac{y^2}{2}+y\bigg]\begin{matrix}
 2 \\
 1
\end{matrix}""=\dfrac{2(2-1)^{3/2}}{3}-\dfrac{(2)^2}{2}+2"

"-(\dfrac{2(1-1)^{3/2}}{3}-\dfrac{(1)^2}{2}+1)=\dfrac{1}{6}({units}^2)"